OpenStudy (anonymous):

Factor completely x^3+5x^2+6x

OpenStudy (triciaal):

x is common then you have a quadratic that can be factored

OpenStudy (triciaal):

do you know how to factor the quadratic?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

@Hotchellerae21

OpenStudy (anonymous):

LOL No.

OpenStudy (anonymous):

Waiting on you to help mee

OpenStudy (anonymous):

the answer is x(x-2)(x-3) first you take out an x then you factor from there

OpenStudy (anonymous):

i mean i hope this helps lol

OpenStudy (triciaal):

x(x^2 + 5x + 6) to factor the quadratic what are the factors of +6 that add up to +5 ? let's use +3 and +2 x^2 + 3x +2x + 6 x(x +2) + 3(x + 2) (x+ 2)(x + 3) back to the original x^3 + 5x^2 + 6x = x(x^2 + 5x +6) = x(x+2)(x+3) any questions?

OpenStudy (triciaal):

sorry, had computer problems it wouldn't post earlier